OpenStudy (anonymous):

Find g(h(x)) let g(x)=-3x^2+7 and h(x)=square root x-2 How do i find this? thanks.. help

OpenStudy (zzr0ck3r):

\(g(h(x)) = g(\color{blue}{\sqrt{x-2}})=-3(\color{blue}{\sqrt{x-2}})^2+7\)

OpenStudy (zzr0ck3r):

simplify...